Assume that a jeweler needs 55 grams of an alloy that is 50% gold. He has available one alloy that is 25% gold, and another that is 70% gold; he will mix some of each.
If x= number of grams of the first alloy (so the number of grams of the second is 55-x), find x.

Answer by jorel1380(3719)

55x.5=27.5
w+z=55
.25w+.7z=27.5
25w+70z=2750
25w+25z=1375
45z=1375
z=30.55555
w=24.44444
You need 24.4444 gms. of 25% gold and 30.555555 gms. of 70% gold for your mix..